OpenStudy (narissa):

evaluate the expression -3x ÷ y for each given x and y value. a) if x =10 and y=-3 b) if x =-10 and y =3 c) if x =-10 and y =-3

OpenStudy (narissa):

@wilsondanielle

OpenStudy (wilsondanielle):

is this the whole question? because it won't equal y in any of those cases.

OpenStudy (narissa):

yes

OpenStudy (wilsondanielle):

start with a, and work through the same process. eg: a) -3x/y = -3*10/-3 = 30/-3 = -10